From year 2001 till 2014, the increase in food prices followed the changes of CPI. The CPI of 2001 was 105.8 and of 2014 125.5. If certain food purchases cost 382 € in year 2014, what was the same food purchases' price in year 2001?

 


1
Expert's answer
2022-02-22T17:11:08-0500

Solution:

The increase in food prices followed the changes of CPI.

It means they are proportional.

Let the same food purchases' price in year 2001 be "x"

"\dfrac{105.8}{125.5}=\dfrac{x}{382}
\\ x=\dfrac{105.8}{125.5}\times382\approx322.036"

So, the same food purchases' price in year 2001 was 322.036 €.